The fundamental vibrational mode for H2 is at 4159/cm. What is the wavelength of the first anti-Stokes transition if the sample is irradiated at a wavelength of 1064nm?
Answer #1

Fundamental vibrational mode for H2 is at, \nu = 4159 / cm

excitation wavlength, \lambda1 = 1064 x 10-7 cm

wavelength of the first anti-Stokes transition = \lambda2

From Raman spectroscopy.

\nu = 1 / \lambda1 - 1 / \lambda2

4159 = (1/1064 - 1/\lambda2) * 10-7

4.159x10-4 = 9.398x10-4 - 1 / \lambda2

1 / \lambda2 = 5.239x10-4

\lambda2 = 1908.76 nm

So,  the wavelength of the first anti-Stokes transition =  1908.58 nm